A company produces steel rods. The lengths of the steel rods are normally distributed with a mean of 170.1-cm and a standard deviation of 1.5-cm. For shipment, 16 steel rods are bundled together. Find the probability that the average length of the rods in a randomly selected bundle is between 169.8-cm and 170-cm. P(169.8-cm < ¯¯¯ X < 170-cm) = Round to 4 decimal places. Answers obtained using exact z-scores or z-scores rounded to 2 decimal places are accepted.
Answer #1

Solution :

= / n = 1.5 / 16 = 0.375

= P[(169.8 - 170.1) / 0.375 < ( - ) / < (170 - 170.1) / 0.375)]

= P(-0.80 < Z < -0.27)

= P(Z < -0.27) - P(Z < -0.80)

= 0.3936 - 0.2119

= 0.1817